OpenStudy (anonymous):

Solve each system by substitution 2x+7y=3 -2x+y=-11 and show work please

OpenStudy (anonymous):

start with the bottom equation -2x + y = -11 solve for y y = -11 + 2x plug this into the first equation wherever you see y 2x+7y=3 2x + 7*(-11+2x) = 3 2x -77 + 14x = 3 16x = 80 x = ? plug the value of x into either equation and solve for y.

OpenStudy (anonymous):

Can you show me how to do y

OpenStudy (anonymous):

it really is as easy as plugging in the value of x into either equation and then just a bit of algebra.

OpenStudy (anonymous):

Just subtract the second equation from the first equation. Your y's will cancel out and then you can solve for x. Then plug your x into either equation and then solve for y

OpenStudy (anonymous):

I mean can you show me using numbers

OpenStudy (anonymous):